Brooks Sporting Inc. is prepared to report the following 2018 income statement (shown in thousands of dollars).

Sales $14,900
Operating costs including depreciation 11,324
EBIT $3,576
Interest 396
EBT $3,180
Taxes (40%) 1,272
Net income $1,908

Prior to reporting this income statement, the company wants to determine its annual dividend. The company has 310,000 shares of common stock outstanding, and its stock trades at $38 per share.

  1. The company had a 35% dividend payout ratio in 2017. If Brooks wants to maintain this payout ratio in 2018, what will be its per-share dividend in 2018? Round your answer to the nearest cent.
    $  

  2. If the company maintains this 35% payout ratio, what will be the current dividend yield on the company's stock? Round your answer to two decimal places.
      %

  3. The company reported net income of $1.7 million in 2017. Assume that the number of shares outstanding has remained constant. What was the company's per-share dividend in 2017? Round your answer to the nearest cent.
    $  

  4. As an alternative to maintaining the same dividend payout ratio, Brooks is considering maintaining the same per-share dividend in 2018 that it paid in 2017. If it chooses this policy, what will be the company's dividend payout ratio in 2018? Round your answer to two decimal places.
      %

  5. Assume that the company is interested in dramatically expanding its operations and that this expansion will require significant amounts of capital. The company would like to avoid transactions costs involved in issuing new equity. Given this scenario, would it make more sense for the company to maintain a constant dividend payout ratio or to maintain the same per-share dividend?
    1. Since the company would like to avoid transactions costs involved in issuing new equity, it would be best for the firm to maintain the same per-share dividend.
    2. Since the company would like to avoid transactions costs involved in issuing new equity, it would be best for the firm to maintain a constant dividend payout ratio.
Answer #1
a.Total Dividend 667.800
Per share Dividend 2.154193548
b.Dividend yield = Expected Dividend/Current Price 5.67%
c.Per share Dividend 1.919354839
d.Payout Ratio 31.184%
e.I.Since the firm would like to avoid transaction costs, best to maintain same per share dividend

1 a. Total Dividend =1908*35% 2 Per share Dividend =B1/310 3 4 b.Dividend yield = Expe=B2/38 =1700*35%/310 6 C.Per share Divi